An 8-m bamboo pole is leaning against the coconut tree. The foot of the pole is 3 meters from the base of the tree. What angle does the pole make with the tree if the pole reaches up to the top of the tree?

Answer by Theo(13342)

the bamboo pole is 8 meters in length.
the distance from the tree is 3 meters.
the tree is perpendicular to the ground (assumed).
you have a right triangle with a hypotenuse of 8 and an adjacent side of 3.
you would use the cosine formula to find the angle.
cosine (angle) = adjacent / hypotenuse = 3/8 = .375
use your calculator to find the angle equals 67.97568716 degrees.
the cosine of 67.97568716 degrees = 3/8.
the arc cosine of 3/8 = 67.97568716 degrees.